# Hook-Lengths, Symplectic/Orthogonal Contents and Amdeberhan’s Conjectures

### v2

@@ Abstract.
The symplectic/orthogonal contents of partitions are related to the dimensions of irreducible representations of symplectic/orthogonal groups. In 2012, motivated by Nekrasov-Okounkov’s hook-length formula and Stanley’s hook-content formula, Amdeberhan proposed several conjectures about infinite product formulas for certain generating functions of hook-lengths and symplectic/orthogonal contents. Some special cases of his conjectures were recently proved by Amdeberhan, Andrews and Ballantine. In this paper, we prove the general cases of Amdeberhan’s conjectures.
@@ 1. Introduction
The study of integer partitions initiated from Euler who showed the generating function of the numbers of partitions as an infinite product formula. Later, various aspects of integer partitions were intensively studied by mathematicians, including their relations to combinatorics, representation theory and mathematical physics. For example, the irreducible representations of symmetric groups, as well as the polynomial representations of general linear groups, can be indexed by partitions. The dimension of representations can be computed by using the hook-lengths of corresponding partitions. In 2006, Nekrasov and Okounkov [14] provided a formula when they studied the Seiberg-Witten theory and random partitions. This formula relates a sum over products of hook-lengths of partitions to the powers of Euler’s infinite product formula. Various new proofs and generalizations of this Nekrasov-Okounkov formula have been explored, as discussed in [8, 9, 10, 17, 21, 23] and references therein. Inspired by Nekrasov-Okounkov’s hook-length formula, and Stanley’s analogous hook-content formula [17], Amdeberhan proposed several conjectures about infinite product formulas for certain generating functions of hook-lengths and symplectic/orthogonal contents [1]. These conjectures have interests in representation theory of symplectic/orthogonal groups (see [6, 12, 20]). Recently, some special cases of Amdeberhan’s conjectures were proved by Amdeberhan, Andrews and Ballantine [2]. In this paper, we will prove the general cases of Amdeberhan’s conjectures. The key point is the use of vertex operators.
@@
Stanley [17], inspired by the Nekrasov-Okounkov formula [14], proposed the hook-content formula, which relates a sum over products of hook-lengths and contents to an infinite product formula. The hook-contents are related to the representation theory of general linear groups. Thus it is natural to consider its symplectic and orthogonal counterparts, which should involve symplectic contents and orthogonal contents. These are exactly the Amdeberhan’s conjectures in [1].
@@
It is known that after taking conjugations of partitions, the symplectic contents and the orthogonal contents are dual to each others. Based on this combinatorial fact, we prove
@@
Conjecture 6.3 (a) in [1] corresponds to the special case $t_{1}=-t_{2}$ of Theorem 1.3. See Example 4.4 for more details. As a consequence of Theorem 1.1 and Theorem 1.3, we have
@@
The cases $t=0$ of Theorem 1.1 and Corollary 1.2, and the cases $t_{1}=-t_{2}=0$ and $t_{1}=-t_{2}=1$ of Theorem 1.3 were recently proved by Amdeberhan, Andrews and Ballantine in [2]. Using the method developed in this paper, we can also provide the following new formula, which involves symplectic contents and ordinary contents at the same time.
@@
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}\prod_{u\in\lambda}\frac{\big{(}t_{1}+c_{sp}(u)\big{)}\big{(}t_{2}+c(u)\big{)}}{h^{2}(u)}=\prod_{n\geq 1}\frac{(1-q^{2})^{\binom{t_{2}}{2}}}{(1-q)^{2t_{1}t_{2}}}.$
@@ 2.1. Schur functions and symplectic Schur functions
The theory of symmetric functions plays important roles in the representation theory. In general, one starts from a set of independent variables $(x_{1},...,x_{n})$ and denote by $\Lambda_{n}(x_{1},...,x_{n})$ the set of symmetric polynomial functions in these variables. It is well-known that the following power sum coordinates
@@
It is more convenient to extend the notion of symmetric functions to functions in infinitely many variables $(\bm{x})=(x_{1},x_{2},...)$. In terms of certain projective limit, one can obtain the ring $\Lambda(\bm{x})$ consisting of symmetric functions in infinitely many variables $(\bm{x})=(x_{1},x_{2},...)$ (for more details, see Chapter I.2 in [13]). Here we are focusing on some concrete symmetric functions. The $n$-th power sum coordinates $p_{n}(\bm{x})$ are still defined by equation (2.1) but they now involve infinitely many variables $(\bm{x})$. General symmetric functions could also be represented as polynomials in terms of these power sum coordinates. For example, the famous Schur functions can be defined in this manner by using generating function of elementary Schur functions and Jacobi-Trudi formula. In this paper, we formally define them as follows. The Schur function labeled by a partition $\lambda=(\lambda_{1},...,\lambda_{l})$ is
@@
A kind of generalization of Schur functions are the skew Schur functions. They are labeled by two partitions. More concretely, for two given partitions $\lambda,\mu$, the skew Schur function labeled by $(\lambda,\mu)$ is defined by
@@
The importance of Schur functions is that they encode information on irreducible representations of symmetric groups, as well as the polynomial representations of general linear groups (see [13, 16]). So do symplectic Schur functions. They encodes information on the irreducible polynomial representations of even symplectic groups (see [20]). For example, the following special evaluation of the symplectic Schur functions (see [4, 6])
@@
$sp_{\lambda}(\tilde{\bm{x}})|_{p_{k}(\tilde{\bm{x}})=2n,\forall k}=sp_{\lambda}(\bm{p})|_{p_{k}=2n}=\prod_{u\in\lambda}\frac{2n+c_{sp}(u)}{h(u)}.$ (2.4)
@@ 2.3. Vertex operator
It is just the homogeneous operator when assigning $\deg q_{k}=k$. The commutation relations between $L_{0}$ and $\Gamma_{\pm}(z)$ are
@@
When regarding $(\bm{p})$ as the power sum coordinates $\big{(}\bm{p}(\bm{y})\big{)}=\big{(}p_{k}(\bm{y})\big{)}$, in the symmetric variables $(\bm{y})=(y_{1},y_{2},...)$, one has
@@
The relation between vertex operators and Schur functions are fundamental (see, for example, equation (A.15) in [15])
@@ 3.1. A partition function of symplectic Schur functions
Theorem 3.1 has the following equivalent formalism expressed in terms of the symmetric variables $(\bm{y})=(y_{i})$.
Corollary 3.2.
@@
Then the right hand side of equation (3.2) is a direct rewriting of equation (3.1) in terms of the symmetric variables $(\bm{y})=(y_{i})$. $\Box$
@@
In this subsection, we prove Theorem 3.1 by showing equation (3.2). First, we need to introduce the following operator
@@
Proof: The first result that $\Gamma_{-}(z)$ and $\exp\Big{(}\sum_{n=1}\frac{1}{n}(\frac{1}{2}\alpha_{-n}^{2}-\frac{1}{2}\alpha_{-2n})\Big{)}$ commute with each other follows from the fact that both of them consist only of $\alpha_{-n}$ for $n>0$. Equation (3.5) is a consequence of the following two commutation relations
$\Gamma_{+}(z)\exp\Big{(}\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-2n}\Big{)}=(1-z^{2})^{-1/2}\cdot\exp\Big{(}\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-2n}\Big{)}\Gamma_{+}(z),$
$\Gamma_{+}(z)\exp\Big{(}-\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-n}^{2}\Big{)}=(1-z^{2})^{1/2}\cdot\exp\Big{(}-\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-n}^{2}\Big{)}\Gamma_{-}(z)^{-1}\Gamma_{+}(z).$
In return, these two equations can be proved by using the Baker-Campbell-Hausdorff formula while the second equation additionally demands the following
$\exp\Big{(}\sum_{n=1}^{\infty}\frac{z^{n}}{n}(\alpha_{n}-\alpha_{-n})\Big{)}=(1-z^{2})^{-1/2}\cdot\Gamma_{-}(z)^{-1}\Gamma_{+}(z),$
which is provable by the Zassenhaus formula. Equation (3.6) can be shown by multiplying both sides of equation (3.5) by $\Gamma_{+}(z)^{-1}$. $\Box$
@@
Another useful operator is
@@
Denote its dual by $F^{*}$. Then we have
@@
With above commutation relations at hand, we are prepared to prove Theorem 3.1.
Proof of Theorem 3.1: Denote by $\mathcal{P}^{\prime}$ the set of partitions whose Frobenius notations are of the form $(m_{1},...,m_{r}|m_{1}+1,...,m_{r}+1)$ for some decreasing non-negative integers $m_{1},...,m_{r}$. For convenience, the empty partition $\emptyset$ is considered in $\mathcal{P}^{\prime}$. The following result was proved by Sundaram in her thesis (see Theorem 12.10 in [18])
@@
where the second equal sign is an application of formula (2.8), which expresses skew Schur function in terms of the vacuum expectation value involving vertex operators.
The following equation is useful (see I.5.Example 9 (a) in [13]), which is in fact our motivation to introduce the operator $G$,
$\sum_{\lambda\in\mathcal{P}^{\prime}}(-1)^{|\lambda|/2}s_{\lambda}\big{(}\bm{p}(\bm{y})\big{)}=\prod_{i<j}(1-y_{i}y_{j}).$
@@
$\prod_{i<j}(1-y_{i}y_{j})=\exp\Big{(}-\sum_{n=1}^{\infty}\frac{1}{2n}\big{(}p_{n}(\bm{y})^{2}-p_{2n}(\bm{y})\big{)}\Big{)}.$
Based on the definition (3.4) of the operator $G$ and the homogeneous operator $L_{0}$, the above two equations can be rewritten as
@@ 3.3. Hook-symplectic/orthogonal contents formulas
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}\prod_{u\in\lambda}\frac{n_{1}+c_{sp}(u)}{h(u)}=\prod_{n=1}^{\infty}\Bigg{(}\frac{1-q^{4n-2}}{1-q^{8n-4}}\cdot\frac{(1-q^{8n-4})^{\binom{n_{1}}{2}}(1-q^{4n-1})^{n_{1}}(1-q^{8n})^{\binom{n_{1}}{2}+n_{1}}}{(1-q^{4n-3})^{n_{1}}(1-q^{8n-6})^{\binom{n_{1}}{2}}(1-q^{8n-2})^{\binom{n_{1}+1}{2}}}\Bigg{)}.$ (3.18)

### v3

@@ Abstract.
The symplectic/orthogonal contents of partitions are related to the dimensions of irreducible representations of symplectic/orthogonal groups. In 2012, motivated by Nekrasov–Okounkov’s hook-length formula and Stanley’s hook-content formula, Amdeberhan proposed several conjectures about infinite product formulas for certain generating functions of hook-lengths and symplectic/orthogonal contents. Some special cases of his conjectures were recently proved by Amdeberhan, Andrews and Ballantine. In this paper, we prove the general cases of Amdeberhan’s conjectures.
footnotetext: 2020 Mathematics Subject Classification: 05A17, 05A15, 11P81.
@@ 1. Introduction
The study of integer partitions was initiated by Euler who showed the generating function of the numbers of partitions as an infinite product formula. Later, mathematicians intensively studied various aspects of integer partitions, including their relations to combinatorics, representation theory, and mathematical physics. For example, the irreducible representations of symmetric groups, as well as the polynomial representations of general linear groups, can be indexed by partitions. The dimension of representations can be computed by using the hook-lengths of corresponding partitions. In 2006, Nekrasov and Okounkov [17] provided a formula when they studied the Seiberg–Witten theory and random partitions. This formula relates a sum over products of hook-lengths of partitions to the powers of Euler’s infinite product formula. Various new proofs and generalizations of this Nekrasov–Okounkov formula have been explored, as discussed in [9, 10, 11, 21, 25, 27] and references therein. Inspired by Nekrasov–Okounkov’s hook-length formula, and Stanley’s analogous hook-content formula [21], Amdeberhan proposed several conjectures about infinite product formulas for certain generating functions of hook-lengths and symplectic/orthogonal contents [1]. These conjectures may have interests in the representation theory of symplectic/orthogonal groups (see [6, 14, 24]) and mathematical physics, especially in integrable systems (see, for examples, [12, 19]). Recently, some special cases of Amdeberhan’s conjectures were proved by Amdeberhan, Andrews and Ballantine [2]. In this paper, we will prove the general cases of Amdeberhan’s conjectures. The key point is the use of vertex operators.
@@
Stanley [21], inspired by the Nekrasov–Okounkov formula [17], proposed the hook-content formula, which relates a sum over products of hook-lengths and contents to a product formula (see Theorem 2.2 in [21]). The hook-contents are related to the representation theory of general linear groups. Thus it is natural to consider its symplectic and orthogonal counterparts, which should involve symplectic contents and orthogonal contents. These are exactly the Amdeberhan’s conjectures in [1].
@@
It is known that after taking conjugations of partitions, the symplectic contents and the orthogonal contents are dual to each other. Based on this combinatorial fact, we prove
@@
Conjecture 6.3 (a) in [1] follows from setting $t_{1}=-t_{2}$ and replacing $t_{2}^{2}$ by $-t$ in Theorem 1.3. See Example 4.4 for more details. As a consequence of Theorem 1.1 and Theorem 1.3, we have
@@
The cases $t=0$ of Theorem 1.1 and Corollary 1.2, and the cases $t_{1}=-t_{2}=0$ and $t_{1}=-t_{2}=1$ of Theorem 1.3 were recently proved by Amdeberhan, Andrews and Ballantine in [2]. Using the method developed in this paper, and Sundaram’s Cauchy identity for symplectic Schur functions (see Theorem 15.1 in [22]), we can also provide the following new formula, which involves symplectic contents and ordinary contents at the same time.
@@
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}\prod_{u\in\lambda}\frac{\big{(}t_{1}+c_{sp}(u)\big{)}\big{(}t_{2}+c(u)\big{)}}{h^{2}(u)}=\frac{(1-q^{2})^{\binom{t_{2}}{2}}}{(1-q)^{t_{1}t_{2}}}.$ (1.8)
@@ 2.1. Schur functions and symplectic Schur functions
The theory of symmetric functions plays an important role in the representation theory. In general, one starts from a set of independent variables $(x_{1},...,x_{n})$ and denote by $\Lambda_{n}(x_{1},...,x_{n})$ the set of symmetric polynomial functions in these variables. It is well known that the following power sum coordinates
@@
It is more convenient to extend the notion of symmetric functions to functions in infinitely many variables $(\bm{x})=(x_{1},x_{2},...)$. In terms of certain projective limit, one can obtain the ring $\Lambda(\bm{x})$ consisting of symmetric functions in infinitely many variables $(\bm{x})=(x_{1},x_{2},...)$ (for more details, see Chapter I.2 in [15]). Here we are focusing on some concrete symmetric functions. The $n$-th power sum coordinates $p_{n}(\bm{x})$ are still defined by equation (2.1), but they now involve infinitely many variables $(\bm{x})$. General symmetric functions could also be represented as polynomials in terms of these power sum coordinates. For example, the famous Schur functions can be defined in this manner by using the generating function of elementary Schur functions and the Jacobi–Trudi formula. In this paper, we formally define them as follows. The Schur function labeled by a partition $\lambda=(\lambda_{1},...,\lambda_{l})$ is
@@
A kind of generalization of Schur functions is the skew Schur functions. They are labeled by two partitions. More concretely, for two given partitions $\lambda,\mu$, the skew Schur function labeled by $(\lambda,\mu)$ is defined by
@@
The importance of Schur functions is that they encode information on irreducible representations of symmetric groups, as well as the polynomial representations of general linear groups (see [15, 20]). So do symplectic Schur functions. They encode information on the irreducible polynomial representations of even symplectic groups (see [24]). For example, the following special evaluation of the symplectic Schur functions (see [4, 6])
@@
$sp_{\lambda}(\tilde{\bm{x}})|_{p_{k}(\tilde{\bm{x}})=2n,\forall k}=sp_{\lambda}(\bm{p})|_{p_{k}=2n,\forall k}=\prod_{u\in\lambda}\frac{2n+c_{sp}(u)}{h(u)}.$ (2.4)
@@ 2.3. Vertex operator
It is the homogeneous operator if we regard $\mathbb{C}[\bm{q}]$ as a graded ring by assigning the degree of $q_{k}$ to be $k$. The commutation relations between $L_{0}$ and $\Gamma_{\pm}(z)$ are
@@
When regarding $(\bm{p})$ as the power sum coordinates $\big{(}\bm{p}(\bm{y})\big{)}=\big{(}p_{k}(\bm{y})\big{)}$, in the variables $(\bm{y})=(y_{1},y_{2},...)$, one has
@@
The relation between vertex operators and Schur functions is fundamental (see, for example, equation (A.15) in [18])
@@ 3.1. A partition function of symplectic Schur functions
Theorem 3.1 has the following equivalent formalism expressed in terms of the variables $(\bm{y})=(y_{i})$.
Lemma 3.2.
@@
Then the right hand side of equation (3.2) is a direct rewriting of equation (3.1) in terms of the variables $(\bm{y})=(y_{i})$. $\Box$
@@
In this subsection, we prove Theorem 3.1 by showing equation (3.2).
For two given operators $A,B$, if we already know their commutation relation, then the following special case of the well-known Baker–Campbell–Hausdorff formula (see, for example, Lemma 5.3 in [16])
$e^{A}Be^{-A}=\sum_{n=0}^{\infty}\frac{1}{n!}ad_{A}^{n}(B),$ (3.4)
where $ad_{A}^{0}(B):=B,ad_{A}^{n}(B)=[A,ad_{A}^{n-1}(B)]$, tells us how to compute the commutation relation involving their exponents. Fortunately, even if there is an infinite sum in this formula, in all of our cases, we have that $ad_{A}^{2}(B)$ is a scalar operator then $ad_{A}^{n}(B)=0$ for all $n>2$. Moreover, if $[A,B]$ is already a scalar operator, then the Zassenhaus formula reads
$e^{A+B}=e^{A}e^{B}e^{-[A,B]/2}.$ (3.5)
These two formulas are the main tools in the following Lemma 3.3.
To prove equation (3.2), we first need to introduce the following operator
@@
The motivation to consider such an operator comes from equations (3.19) and (3.20), which will be exhibited later. Below, we first study the commutation relations with respective to $\exp(G)$.
@@
Proof: First, since both of the operators $\Gamma_{-}(z)$ and $\exp\Big{(}\sum_{n=1}\frac{1}{n}(\frac{1}{2}\alpha_{-n}^{2}-\frac{1}{2}\alpha_{-2n})\Big{)}$ consist only of $\alpha_{-n}$ for $n>0$, they commute with each other. Next, one can use the Heisenberg relations $[\alpha_{m},\alpha_{n}]=m\delta_{m+n,0}$ to obtain
$\big{[}\sum_{n=1}^{\infty}\frac{z^{n}\alpha_{n}}{n},\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-2n}\big{]}=\sum_{n=1}^{\infty}\frac{z^{2n}}{2n}=-\frac{1}{2}\ln(1-z^{2}).$
Then, one can directly apply the Baker–Campbell–Hausdorff formula (3.4) to obtain the following commutation relation
$\Gamma_{+}(z)\exp\Big{(}\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-2n}\Big{)}=(1-z^{2})^{-1/2}\cdot\exp\Big{(}\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-2n}\Big{)}\Gamma_{+}(z).$ (3.9)
By the similar method, one can derive
$\Gamma_{+}(z)\exp\Big{(}-\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-n}^{2}\Big{)}=(1-z^{2})^{1/2}\cdot\exp\Big{(}-\sum_{n=1}^{\infty}\frac{1}{2n}\alpha_{-n}^{2}\Big{)}\Gamma_{-}(z)^{-1}\Gamma_{+}(z)$ (3.10)
using the Baker–Campbell–Hausdorff formula (3.4). It additionally demands the following
$\exp\Big{(}\sum_{n=1}^{\infty}\frac{z^{n}}{n}(\alpha_{n}-\alpha_{-n})\Big{)}=(1-z^{2})^{-1/2}\cdot\Gamma_{-}(z)^{-1}\Gamma_{+}(z),$ (3.11)
which is provable by the Zassenhaus formula (3.5). Then, equation (3.7) is a consequence of the above two commutation relations (3.9) and (3.10). Equation (3.8) can be shown by multiplying both sides of equation (3.7) by $\Gamma_{+}(z)^{-1}$. $\Box$
@@
Motivated from the I.5.Example 4 in [15] (see also equation (5.2) in [28] for the following equivalent form),
$\sum_{\lambda\in\mathcal{P}}s_{\lambda}(\bm{p})=\exp\Big{(}\sum_{n=1}^{\infty}\frac{1}{n}\big{(}p_{n}+\frac{1}{2}p_{n}^{2}-\frac{1}{2}p_{2n}\big{)}\Big{)},$
it is beneficial to introduce another useful operator as
@@
Denote its dual by $F^{*}:=\sum_{n=1}^{\infty}\frac{1}{n}(\alpha_{n}+\frac{1}{2}\alpha_{n}^{2}-\frac{1}{2}\alpha_{2n})$. Then we have
@@
With the above commutation relations at hand, we are prepared to prove Theorem 3.1.
Proof of Theorem 3.1: Denote by $\mathcal{P}^{\prime}$ the set of partitions whose Frobenius notations are of the form $(m_{1},...,m_{r}|m_{1}+1,...,m_{r}+1)$ for some decreasing non-negative integers $m_{1},...,m_{r}$. Actually, after taking the conjugation, this kind of partition gives the doubled distinct partition studied in [8, 14]. For convenience, the empty partition $\emptyset$ is considered in $\mathcal{P}^{\prime}$. The following result was proved by Sundaram in her thesis (see Theorem 12.10 in [22])
@@
where the second equality sign is an application of formula (2.8), which expresses the skew Schur function in terms of the vacuum expectation value involving vertex operators.
The following equation is useful (see I.5.Example 9 (a) in [15]),
$\sum_{\lambda\in\mathcal{P}^{\prime}}(-1)^{|\lambda|/2}s_{\lambda}\big{(}\bm{p}(\bm{y})\big{)}=\prod_{i<j}(1-y_{i}y_{j}).$ (3.19)
@@
$\prod_{i<j}(1-y_{i}y_{j})=\exp\Big{(}-\sum_{n=1}^{\infty}\frac{1}{2n}\big{(}p_{n}(\bm{y})^{2}-p_{2n}(\bm{y})\big{)}\Big{)}.$ (3.20)
The operator $G$, defined in equation (3.6), is indeed the operator by multiplying the function above. Moreover, together with the homogeneous operator $L_{0}$, the above two equations (3.19) and (3.20) can be rewritten as
@@ 3.3. Hook-symplectic/orthogonal contents formulas
$\begin{split}\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}\prod_{u\in\lambda}&\frac{n_{1}+c_{sp}(u)}{h(u)}\\ &=\prod_{n=1}^{\infty}\Bigg{(}\frac{1-q^{4n-2}}{1-q^{8n-4}}\cdot\frac{(1-q^{8n-4})^{\binom{n_{1}}{2}}(1-q^{4n-1})^{n_{1}}(1-q^{8n})^{\binom{n_{1}}{2}+n_{1}}}{(1-q^{4n-3})^{n_{1}}(1-q^{8n-6})^{\binom{n_{1}}{2}}(1-q^{8n-2})^{\binom{n_{1}+1}{2}}}\Bigg{)}.\end{split}$ (3.25)